Your college newspaper, The Collegiate Investigator, has fixed production costs of $76 per edition and marginal printing and distribution costs of 41¢/copy. The Collegiate Investigator sells for 51¢/copy.

(a) Write down the associated cost function C(x) in dollars.
C(x) =

Write down the revenue function R(x) in dollars.
R(x) =

Write down the profit function P(x) in dollars.
P(x) =

(b) What profit (or loss) results from the sale of 500 copies of The Collegiate Investigator?
$
(c) How many copies should be sold in order to break even?
copies

Answer by xinxin(76)

a). C(x) = 76 + 0.41x
R(x) = 0.51x
P(x) = R(x) - C(x) = 0.1x -76
b). P(500) = 0.1 * 500 -76 = -26 (loss)
c). 0.1x - 76 =0
x = 760